Six people can paint 4 walls in 41 minutes. How many minutes will it take for 10 people to paint 10 walls? Round to the nearest minute.
Answer by srinivas.g(540)

6 people can paint 4 walls in 41 minutes
In one minute 6 people can paint 4/41 walls
In one minute 1 person can paint 4/(41*6) walls
In one minute 10 people can paint 4 *10/(41*6) = 40 /246 wall
= 20/123 wall
to piant one full wall , 10 people take = 123/20 minutes
to paint 10 walss , 10 people take = 123/ 20 *10
= 123/2
= 61.5 minutes
answer = 61.5 minutes